OpenStudy (anonymous):

Which equation has a solution of n equals negative one ? A)One half n minus six equals negative seven B)Negative six n plus eight equals negative two C)Two times n minus eight equals negative eighteen D)Two n plus twelve minus three n equals fifteen

OpenStudy (texaschic101):

1/2n - 6 = -7 1/2n = -7 + 6 1/2n = -1 n = -1/(1/2) n = -1 * 2 n = -2 not A. -6n + 8 = -2 -6n = -8 - 2 -6n = - 10 n = 5/3 not B 2n - 8 = -18 or 2(n - 8) = -18 2n = -18 + 8 n - 8 = -9 2n = -10 n = -9 + 8 n = -5 n = -1 not C This is it 2n + 12 - 3n = 15 -n + 12 = 15 -n = 15 - 12 -n = 3 n = -3 not D At first I didn't think the answer was here. The answer is C, but only when parenthesis are inserted in the equation.